ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 174, 2009

Adan-Bante, E. Restriction of characters and products of characters  221

Alon, N. and E. Lubetzky Poisson approximation for nonbacktracking random walks

227

Alon, N., B. Bukh and B. Sudakov Discrete Kakeya-type problems and small bases

285

Alonso Álvarez, J. N., J. M. Fernández Vilaboa, R. González Rodríguez and M. P. López A characterization of projective weak Galois extensions

161

Bazzoni, S. and D. Herbera Cotorsion pairs generated by modules of bounded projective dimension

119

Beals, R. M., J. L. Wetherell and M. E. Zieve Polynomials with a common composite

93

Bergelson, V. and I. Z. Ruzsa Sumsets in difference sets  1

Boshernitzan, M. and D. Damanik The repetition property for sequences on tori generated by polynomials or skew-shifts

189

Bukh, B. See N. Alon Casolo, C. and S. Dolfi Products of primes in conjugacy class sizes

and irreducible character degrees

403

Chiang, R. and E. Z. Xia Strong Lefschetz property under reduction  369
Collins, D. J. and A. Juhász Engel Groups II  47
Collins, D. J. and A. Juhász Engel groups III  73

Damanik, D. See M. Boshernitzan Dolfi, S. See C. Casolo Enochs, E. E. and H. Holm Cotorsion pairs associated with Aus-

lander categories

253

Fernández Vilaboa, J. M. See J. N. Alonso Álvarez González Rodríguez, R. See J. N. Alonso Álvarez Herbera, D. See S. Bazzoni Holm, H. See E. E. Enochs Juhász, A. See D. J. Collins Kahane, J.-P. and Y. Katznelson Restrictions of continuous func-

tions

269

Katznelson, Y. See J.-P. Kahane Kiessling, J. Classification of certain cellular classes of chain com-

plexes

179

López, M. P. See J. N. Alonso Álvarez Li, J., H. Liu and F. Qiu Exact and explicit analytic solutions of an

extended Jabotinsky functional differential equation

29

Liu, H. See J. Li Lubetzky, E. See N. Alon Moravec, P. Groups of prime power order and their nonabelian ten-

sor squares

19

Nevo, E. and U. Wagner On the embeddability of skeleta of spheres  381

Qiu, F. See J. Li Roychowdhury, M. K. and D. J. Rudolph Any two irreducible

Markov chains are finitarily orbit equivalent

349

Rudolph, D. J. See M. K. Roychowdhury Ruzsa, I. Z. See V. Bergelson Suciu, L. Maximum A-isometric part of an A-contraction and appli-

cations

419

Sudakov, B. See N. Alon Wagner, U. See E. Nevo

Wen, Z.-Y. and L.-F. Xi The geometry of Whitney’s critical sets  303

Wetherell, J. L. See R. M. Beals Xia, E. Z. See R. Chiang Xi, L.-F. See Z.-Y. Wen

Yadin, A. When do random subsets decompose a finite group?  203

Zieve, M. E. See R. M. Beals